# Diagnostic Reasoning in Natural Language: Computational Model and Application

### v1

@@ Abstract
Diagnostic reasoning is a key component of expert work in many domains. It is a hard, time-consuming activity that requires expertise, and AI research has investigated the ways automated systems can support this process. Yet, due to the complexity of natural language, the applications of AI for diagnostic reasoning to language-related tasks are lacking. To close this gap, we investigate diagnostic abductive reasoning (DAR) in the context of language-grounded tasks (NL-DAR). We propose a novel modeling framework for NL-DAR based on Pearl’s structural causal models and instantiate it in a comprehensive study of scientific paper assessment in the biomedical domain. We use the resulting dataset to investigate the human decision-making process in NL-DAR and determine the potential of LLMs to support structured decision-making over text. Our framework, open resources and tools lay the groundwork for the empirical study of collaborative diagnostic reasoning in the age of LLMs, in the scholarly domain and beyond. [^1] Code and data will be publicly released upon acceptance.
@@ 1 Introduction
Diagnostic abductive reasoning (DAR) is an important part of human decision-making in expert domains. Given observations about a system, DAR seeks to find the most plausible explanation for the observed behavior in terms of the system’s hidden states Reiter (1987). Applications of DAR are manifold: doctors use diagnostic reasoning to determine the disease based on symptoms and measurements, and programmers diagnose code to determine the faulty software component that leads to the observed runtime behavior Magnani (2023). Diagnostic reasoning is a hard expert task, and a long line of research has explored the use of AI to expedite this process (Peng and Reggia, 1990; Richens et al., 2020, i.a.). Recent applications of AI to support DAR that decompose problems of the physical world into step-wise diagnostic graphs, such as in the medical domain, show great promise Richens and Buchard (2022).
Figure 1: We use NL-DAR framework to decompose paper assessment into a set of causally connected steps, incl. reading the paper, extracting the evidence, reasoning over the evidence, and arriving at a verdict.
Assessing documents written in natural language is a key part of day-to-day knowledge work, and it involves a lot of diagnostic thinking. From essay grading to fact-checking to scholarly peer review, experts reason about a document based on specific observations (e.g., "Does the news report cite its sources?", "Does the paper report participant details?") and arrive at a conclusion ("the article might be unreliable", "the results might be biased"). Yet, for most textual assessment tasks, only the outcome (e.g., a factuality label or the review report) and not its underlying reasoning process are recorded. This leads to data scarcity that hinders the training and evaluation of natural language processing models using the existing DAR methodology. The potential of natural language DAR (NL-DAR) thus remains under-explored.
Recent breakthroughs in large language models (LLMs) create new opportunities for applying AI to diagnostic reasoning grounded in natural language. In particular, the ability to perform tasks in a zero-shot fashion and to reason over the inputs Brown et al. (2020); Kojima et al. (2022) can alleviate the current methodological hurdles that prevent the application of DAR to natural language problems. Explicit reasoning frameworks, in turn, have the potential to enhance the interpretability and transparency of LLM applications to complex diagnostic tasks, and to steer human-AI collaboration. Thus, investigating LLMs in the context of diagnostic reasoning bears great promise for both lines of research.
Yet, open questions remain. Diagnostic reasoning commonly involves multiple interconnected reasoning steps. (RQ1) How to represent the diagnostic reasoning processes for NL-DAR problems? To approach this, we introduce workflows – an application of structural causal models Pearl (1995) that decompose diagnostic reasoning into a step-wise reasoning graph starting with the text observations and leading towards an outcome. (RQ2) How to devise a workflow for a particular NL-DAR application? We introduce and test a three-step process where we construct workflows based on expert interactions and use the resulting causal models to empirically study how humans perform the task. (RQ3) How can state-of-the-art large language models support NL-DAR? To answer, we perform extensive experiments with four state-of-the-art LLMs and report on the ability of LLMs to perform individual diagnostic reasoning steps as part of an NL-DAR workflow.
Diagnostic reasoning with natural language covers a wide spectrum of problems. To illustrate the application of our framework, we report on a large-scale study in biomedical paper assessment – a task closely related to the emerging line of research in AI-assisted peer review. We enroll over $40$ researchers and students in the biomedical field and develop a practical annotation interface based on the CARE platform Zyska et al. (2023) that allows users to manually perform workflow-based assessment of texts. The resulting large-scale dataset provides an empirical basis for the study of human diagnostic reasoning and LLM assistance in the domain of scholarly peer review, and the toolkit enables applying our proposed framework to other NL-DAR problems. In total, this work contributes:
  • a formal model for natural language abductive diagnostic reasoning tasks rooted in causality theory (RQ1) and a framework for applying this model to specific NL-DAR tasks (RQ2);
@@
  • the first expert-generated multi-modal diagnostic reasoning dataset in biomedical paper assessment, comprising of more than $4$k text-based reasoning steps on top of $22$ biomedical papers;
@@
  • insights into human influence factors during paper assessment (RQ2) and the potential of LLM assistants in diagnostic abductive reasoning for paper assessment (RQ3).
@@
In our experiments, we find that (RQ2) differences in prior knowledge are a central source of disagreement among humans and that a paper’s writing style has a major impact on the overall assessment. We further find that (RQ3) LLMs suffer from error propagation in the reasoning graph but that human oversight effectively mitigates this issue, showing the overall promise of the proposed approach. Our study bridges the gap between the research in causality, human-AI collaboration, and LLMs, and lays the groundwork for applying LLMs to expert tasks that involve complex diagnostic reasoning grounded in natural language.
@@ 2 Related Work
We contribute to abductive reasoning research in NLP Bhagavatula et al. (2020). Prior studies primarily focus on abduction based on formal logic converted into natural language (e.g. Young et al., 2022). In contrast, we take an application-driven approach, which involves complex, subjective expert reasoning on a real-world task. Similar to our study, Dalal et al. (2024) evaluate LLMs’ ability to infer the best explanation for an observation and provide its reasoning. Unlike this work, we impose a reasoning structure on the model to ensure high-quality explanations. Our work is related to research on explainability in NLP (e.g. Wiegreffe and Marasovic, 2021; Camburu et al., 2018). In particular, Dalvi et al. (2021) test models on constructing reasoning paths over short facts (entailment trees) as explanations for science questions. Our framework generalizes entailment trees to directed graphs that serve as explanations. Our work connects NLP with research on AI for DAR. Existing works in this area focus on medical Richens and Buchard (2022) and engineering problems Koitz-Hristov and Wotawa (2018) that rely primarily on bipartite diagnosis graphs. Our framework for DAR in natural language considers graphs of more than two layers; it adds natural language variables to the diagnosis graph and is the first to use LLMs to model the relations between diagnostic variables. Finally, we contribute to NLP for peer review support Kuznetsov et al. (2024). We complement prior work on argumentation and intentions in review reports Hua et al. (2019); Cheng et al. (2020); Kuznetsov et al. (2022); Dycke et al. (2023) by the first empirical analysis of the reasoning that underlies review writing. D’Arcy et al. (2024) and Chamoun et al. (2024) propose multi-agent LLM architectures to generate feedback on scientific papers. In contrast, we test LLMs on the reasoning steps leading towards the assessment of a paper.
@@ 3.1 Background: Causal Models
Causality provides a formal way to reason about cause-and-effect relationships by specifying how active changes to one variable (interventions) change the likelihood of another. One widely used type of causal models, adopted in this work, are Structural Causal Models (SCMs, Pearl (1995)). Following the definition by Pearl (2000):
@@
An SCM $M$ is $(U,V,F,P_{M})$, where
  1. i.
    $U$ is a set of exogenous variables determined by factors outside of $M$. $P_{M}$ is a probability function over the domain of $U$.
  2. ii.
  3. iii.
    $F=\{f_{1},...,f_{n}\}$ is a set of functions on the domains of $U_{i}\cup\textit{pa}_{i}$ to $V_{i}$, where $\textit{pa}_{i}\subseteq V\setminus V_{i}$, and $U_{i}\subseteq U$ and the entire set $F$ forms a mapping from $U$ to $V$.
Causal Graphs
Interventions and Counterfactuals
We notate interventions using the $do$-operator. For an intervention $do(X=x)$, we alter the SCM $M$ by replacing all occurrences of variable $X$ by $x$ in $F$. We denote the altered model as $M_{X=x}$. Counterfactuals pose what-if questions on the SCM. For a query "What value had $Y$ if $X=x$ under the circumstances $u$?" we imagine the causal effect of $X$ on $Y$ for a concrete instance counter to the observed facts. The counterfactual distribution is
$P_{M_{X=x}}(Y=y)=\sum_{\{u|u\in\widehat{U}\land Y_{x}(u)=y\}}P(u)$
where $Y_{x}(u)$ denotes the potential response of $Y$ under $M_{X=x}$. Finally, the average causal effect (ACE) of an intervention on a binary variable $X$ is the difference of expected values of $Y$ with $\textit{ACE}(X,Y)=\mathbb{E}[Y_{x=1}]-\mathbb{E}[Y_{x=0}]$. We use ACE and counterfactuals during analysis in Section 4.3.
3.2 Natural Language Diagnostic Reasoning
At the heart of diagnostic abductive reasoning lies a system model that describes the relations between components and the observable behavior Peng and Reggia (1990). We now formalize an NL-DAR problem using SCMs to specify the system model. [^2] Various formalisms have been proposed for DAR, including fuzzy logic, and probabilistic graphical models Peng and Reggia (1990); Richens et al. (2020). SCMs are strictly more powerful than these formalisms Pearl (2000).
Definition 2
A natural language diagnostic abductive reasoning problem is given by an SCM $M=(U,V,F,P_{M})$, the workflow, with
@@ 3.2 Natural Language Diagnostic Reasoning
    final verdicts on the input with $T\subseteq V$ and $T\neq\emptyset$ analogous to the system behavior.
@@
    components $C\subset V$ with $C\cap I=\emptyset$, $C\cap T=\emptyset$ and $C\neq\emptyset$ forming the explanation, i.e. the steps of the workflow.
@@
All inputs $i\in I$ are roots of the causal graph, i.e., $\textit{pa}_{i}=\emptyset$, and form input to the structural equation of at least one component, i.e., $\forall i\in I:\exists c\in C:i\in\textit{pa}_{c}$. The final verdicts have no children $\forall t\in T:\forall v\in V:t\notin\textit{pa}_{v}$. Background variables determine the level of subjectivity and noise in the task. Solving an NL-DAR problem means inferring the most likely values for all $c\in C$ given $I$ and $T$.
Less formally, a workflow describes a reasoning graph from an input text to the verdict, and the diagnosis is the answers to all intermediate steps explaining the verdict and depending on the textual input. Notably, Definition 2 deviates from DAR problems in the physical world, as the observation (verdict) does not exist on its own but arises as a result of the interpretation of the text by a reader.
3.3 Approaching an NL-DAR Problem
Our definition of NL-DAR implies a three-step approach to problems of this class.
Step 1. Designing the SCM
@@ 3.3 Approaching an NL-DAR Problem
First, the endogenous variables that might impact the final verdicts must be defined. Following that, the SCM’s structure is inferred from observational data through causal discovery Verma and Pearl (1990) or determined through interaction with domain experts. The steps of the workflow are modeled as boolean decisions, numerical scores, or natural language answers [^3] Causality-aware text representation are a hard research problem without a general solution Veitch et al. (2020); Feder et al. (2022). In our case study (Section 4), we combine boolean and natural language answers.. This process yields the graph structure and variables of the workflow.
Step 2. Populating the SCM
@@
Next, structural equations are estimated empirically. To do this, human annotators are guided through the workflow, starting with the input document and moving toward the output verdicts. The intermediate steps serve as the individual explanations of the human’s final verdicts. Varying human experiences and biases act as background factors $U$. If a diagnostic task is subjective, background factors induce high variability in the answers. To measure the level of noise, each step and input is annotated redundantly. Based on this data, one can estimate the structural equations using machine learning models trained on the input-output data for each workflow step.
Step 3. Analysis and automation
Once the underlying SCM is defined and populated, one can analyze it to empirically describe the human diagnostic reasoning process through causal queries on the model. The SCM can be further used as the backbone to solve new instances of specific NL-DAR problems automatically or via human-AI collaboration. To solve new instances of the NL-DAR problem, causal inference is applied, providing new inputs to the trained machine learning models of the structural equations. As we show in Section 5, LLMs naturally combine with the causal graph to solve NL-DAR problems with natural language reasoning steps, devising the SCM as a probabilistic LLM program Dohan et al. (2022).
4 Paper Assessment as NL-DAR
Figure 2: Paper assessment workflow (excerpt). Information from the paper is acquired and aggregated in intermediate reasoning steps. The overall assessment lies at the end of the graph. Each node is associated with a prompt, description, example, and output format. The arrows are causal dependencies reading as therefore.
The growing publication rates across sciences put a strain on academic quality control mechanisms like peer review, which could greatly benefit from AI assistance Kuznetsov et al. (2024). We now report on the application of the proposed framework to biomedical paper assessment cast as an NL-DAR problem. Following Definition 2, our workflow describes the process where humans extract bits of information from the paper text and come to an overall judgment (the verdict) based on a mental model of what constitutes a high-quality paper.

### v2

# STRICTA: Structured Reasoning in Critical Text Assessment for Peer Review and Beyond

@@ Abstract
Critical text assessment is at the core of many expert activities, such as fact-checking, peer review, and essay grading. Yet, existing work treats critical text assessment as a black box problem, limiting interpretability and human-AI collaboration. To close this gap, we introduce Structured Reasoning In Critical Text Assessment (STRICTA), a novel specification framework to model text assessment as an explicit, step-wise reasoning process. STRICTA breaks down the assessment into a graph of interconnected reasoning steps drawing on causality theory Pearl (1995). This graph is populated based on expert interaction data and used to study the assessment process and facilitate human-AI collaboration. We formally define STRICTA  and apply it in a study on biomedical paper assessment, resulting in a dataset of over 4000 reasoning steps from roughly 40 biomedical experts on more than 20 papers. We use this dataset to empirically study expert reasoning in critical text assessment, and investigate if LLMs are able to imitate and support experts within these workflows. The resulting tools and datasets pave the way for studying collaborative expert-AI reasoning in text assessment, in peer review and beyond. [^1] Code and data are publicly available at https://github.com/UKPLab/acl2025-stricta
@@ 1 Introduction
During critical text assessment, an expert evaluates a document and arrives at a verdict about the document’s quality. It is a key task in many expert domains, and often serves as a cornerstone for quality assurance in the respective area Lewandowsky et al. (2020); Johnson et al. (2018); Royce Sadler (2012). Critical text assessment is challenging: the expert must analyze the document from multiple perspectives, integrate background knowledge, and form coherent reasoning to arrive at a judgment. To support the experts, prior work has explored various approaches to automated text assessment in specific domains, such as document quality classification Shen et al. (2019); Maillette de Buy Wenniger et al. (2020), peer review report generation D’Arcy et al. (2024); Yuan et al. (2022); Du et al. (2024); Yu et al. (2024), or automatic essay grading Misgna et al. (2025); Han et al. (2024). Yet, prior work largely treats critical text assessment as a black box, leaving a crucial question unanswered: How do the experts arrive at their judgement? Answering this question is a key prerequisite for transparent and reliable human-AI collaboration in critical text assessment. It would enable comparative analysis of human decision-making, provide scaffolding for fine-grained evaluation of AI assistance, and facilitate human-AI collaboration via enhanced interpretability and human oversight Luo et al. (2024); Ferdaus et al. (2024). Yet, the lack of formal models and datasets that capture expert reasoning during critical text assessment prevents progress in this area. Our work aims to address this gap.
Figure 1: Structured Reasoning in Critical Text Assessment (STRICTA) represents critical text assessment as a reasoning process decomposed into causally linked steps. Based on the observed text, the expert produces a verdict by filling-in each step.
To formalize expert reasoning, we introduce STRICTA: a specification framework that represents critical text assessment as a step-wise structured reasoning graph (workflow), capturing how different aspects of a document contribute to the final evaluation (Figure 1). We propose a three-step process for constructing and populating workflows based on expert interactions. Since many text assessment tasks are partially subjective and allow for multiple valid interpretations, we explore the use of structural causal models (SCMs) Pearl (2000) to explicitly account for this variance during reasoning. We also investigate how workflows can be used to evaluate LLMs’ ability to imitate human reasoning during critical text assessment, casting it as an abductive reasoning problem.
The assessment of scientific papers underlying the peer review process is arguably one of the most challenging document assessment tasks requiring complex reasoning. To demonstrate the use of STRICTA in practice, we therefore focus on paper assessment – a critical text assessment task closely related to the emerging line of work on AI-assisted peer review Shah (2022); Kuznetsov et al. (2024); Goldberg et al. (2024). Specifically, we conduct a large-scale study in biomedical paper assessment. We construct reasoning workflows based on expert interviews, and engage over $40$ biomedical researchers and students to collect workflow-based assessments for more than $20$ papers. With over $4000$ explicit reasoning steps, this unique dataset provides an empirical foundation for studying how experts reason during text assessment, and how modern LLMs can assist them in this task. Overall, we contribute:
  • a model for STRICTA tasks rooted in causality, and a framework for its application;
@@
  • a unique expert-generated multi-modal structured reasoning dataset in biomedical paper assessment;
@@
  • insights into human influence factors during paper assessment and the potential of LLM assistants in STRICTA  applied to paper assessment.
@@
Our analysis reveals that differences in prior knowledge are a key source of disagreement among experts, while a paper’s writing style has a major impact on the overall assessment. Experiments with four state-of-the-art LLMs demonstrate that LLMs are prone to error propagation, but human oversight effectively mitigates this issue. Our study connects research on causality, expert-AI collaboration, and LLMs and paves the path towards transparent and reliable AI assistance in expert domains.
@@ 2 Related Work
Our work offers a novel perspective on critical text assessment with a practical focus on research paper assessment. It directly relates to research on peer review generation and other document quality assessment domains such as fact-checking. The proposed approach is also inspired by research on abductive and diagnostic reasoning and relates to causality in NLP, which we briefly cover in Appendix A.1. Existing work on text assessment trains models to predict the quality of documents such as scientific papers or Wikipedia articles using external quality labels (e.g., citation counts) Shen et al. (2019); Maillette de Buy Wenniger et al. (2020); Wenniger et al. (2023). However, these methods rely on proxy indicators without human validation and treat document assessment as a black-box problem. Our work complements these approaches by focusing on human assessments and the reasoning process to enhance interpretability. Automatic review generation aims to produce full peer review reports for papers using specialized model architectures D’Arcy et al. (2024); Yuan et al. (2022), testing LLMs out-of-the-box Du et al. (2024); Tyser et al. (2024); Liu and Shah (2023); Yu et al. (2024), or providing isolated feedback on paragraphs Chamoun et al. (2024). While these methods offer valuable insights into downstream performance in peer review, they neglect the fine-grained decision-making factors in humans and LLMs that shape the complex assessment. Our study focuses on the underlying structured reasoning process rather than generating a review report, enabling a deeper analysis. Another related area of research is explanations in fact-checking, focused on assessing the veracity of text Guo et al. (2022). Prior studies have generated structured explanations for final verdicts as rule chains Gad-Elrab et al. (2019), argumentation graphs Si et al. (2024), multi-hop graphs Jiang et al. (2020), or verification programs Pan et al. (2023), focusing on short factual claims and often ignoring ambiguity Glockner et al. (2024). In contrast, our approach explicitly models the uncertainties induced by subjectivity and ambiguity, explores diverse decision-making factors in humans, and focuses on whole documents.
@@ 3.1 Background: Causal Models
As causal models are not ubiquitous in NLP, we provide a short introduction. Causality offers a formal framework for analyzing cause-and-effect relationships, i.e., the impact of active changes to one variable (interventions) on the likelihood of another. This work adopts Structural Causal Models (SCMs, Pearl (2000)), a widely used class of causal models. For brevity, we provide a high-level overview of SCMs and their associated concepts. Appendix A.2.1 provides the formal definitions.
Following Pearl (2000), an SCM $\mathcal{M}$ is a tuple $(U,V,F,P_{\mathcal{M}})$ where $U$ are the background variables distributed by $P_{\mathcal{M}}$, $V$ are the modeled variables, and $F$ is the set of structural equations. Each modeled variable is given by a structural equation $v_{i}=f_{i}(\textit{pa}_{i},u_{i})$ where $\textit{pa}_{i}$ is the set of parent variables from $V$ and $U$. The background factors $U$ add randomness to the otherwise deterministic structural equations. We represent SCMs by a causal graph. An edge from $v_{1}$ to $v_{2}$ indicates that $v_{1}$ is input to the structural equation of $v_{2}$ (the child). For an intervention $do(X=x)$, we alter the SCM $\mathcal{M}$ by replacing $X$ by $x$ in $F$ resulting in $\mathcal{M}_{X=x}$. Counterfactuals pose what-if questions on the SCM. For a query, "What value had $Y$ if $X=x$ under the circumstances $u$?" we imagine the effect of $X=x$ on $Y$ for a concrete instance counter to the observed facts. The average causal effect (ACE) of an intervention on a binary variable $X$ is the difference of expected values for $Y$ with $\textit{ACE}(X,Y)=\mathbb{E}[Y_{x=1}]-\mathbb{E}[Y_{x=0}]$ for $Y_{x}(u)$ the potential response of $Y$ under $\mathcal{M}_{X=x}$. We use ACE and counterfactuals during analysis (Sec. 4.3).
3.2 Structured Reasoning in Critical Text Assessment (STRICTA)
We define our Structured Reasoning in Critical Text Assessment framework to model a wide range of document assessment tasks with minimal assumptions. This framework assumes an expert extracts key information from the document as input for a series of interconnected reasoning steps. These steps culminate in a final verdict on the document’s quality and provide explanations linking the text to the outcome. The reasoning steps represent the typically unobserved mental processes that influence the assessment, with specific steps and connections varying based on the problem domain. We formalize the STRICTA problem family using SCMs, as they directly represent causal reasoning structures and facilitate the analysis of explanatory behavior in humans and AIs. Unlike purely statistical models, SCMs allow reasoning about causal rather than merely associational relationships between reasoning steps. Thus, causal modeling is indispensable for studying decision factors during text assessment. [^2] In Section 4.3, we make use of these capabilities posing complex causal queries on the constructed SCM for paper assessment. Appendix A.2.2 discusses alternative formalisms.
@@ 3.2 Structured Reasoning in Critical Text Assessment (STRICTA)
A structured reasoning for critical text assessment problem is given by an SCM $\mathcal{M}=(U,V,F,P_{\mathcal{M}})$, the workflow, with
@@
    final verdicts on the document with $T\subseteq V$ and $T\neq\emptyset$.
@@
    reasoning components $C\subset V$ with $C\cap I=\emptyset$, $C\cap T=\emptyset$ and $C\neq\emptyset$ forming the steps of the workflow.
@@
Furthermore, all inputs are roots of the workflow graph and there is a path that connects each step with at least one input. The final verdicts are terminal nodes in the graph. The background variables determine the level of subjectivity and noise in the task. Solving a STRICTA problem means finding the most likely values for all unobserved $E_{\text{hidden}}\subset C\cup T$ given $I$ and partially observed values $E_{\text{observed}}\subset C\cup T$ with $E_{\text{hidden}}\cap E_{\text{observed}}=\emptyset$.
Informally, a workflow describes a reasoning graph that connects an input text to a final verdict. While the assessment structure remains fixed, the answers for each step vary by instance. The reasoning task depends on the observed reasoning steps. In Section 5, we present a challenging abductive reasoning problem, providing the LLM only with the paper text and final verdict, and tasking it with reconstructing the intermediate reasoning steps.
Notably, we assume that all problem instances in a STRICTA problem class share the same reasoning structure, with variation arising only in the graph’s variables. This contrasts with prior work that constructs reasoning structures for each instance individually (e.g. in fact checking Pan et al., 2023). A fixed reasoning structure, where each step has a well-defined, context-independent meaning, enables quantitative comparison of human reasoning factors and fine-grained evaluation of automatic methods across instances.
3.3 Instantiating a STRICTA Problem
STRICTA represents a family of document assessment problems. Addressing a specific problem, such as biomedical paper assessment, involves the following steps:
@@ 3.3 Instantiating a STRICTA Problem
Step 1. Designing the SCM’s Structure: First, we specify the steps and causal links of the workflow graph. Ultimately, this means determining the influence factors on the final verdict based on the underlying document. This can either be done automatically through causal discovery Verma and Pearl (1990) or through interaction with domain experts. The steps of the workflow can be modeled as boolean decisions, numerical scores, or text answers. [^3] Causality-aware text representation is a hard research problem without a general solution Veitch et al. (2020); Feder et al. (2022). In our case study (Section 4), we combine boolean and natural language answers. This process yields the graph structure and variables of the workflow.
@@
Step 2. Populating the SCM with Data: Next, we determine the relationships between the workflow steps by estimating the SCM’s structural equations. To do this, human annotators follow the workflow to assess the paper, answering each step after reading the document and forming a verdict. By enforcing the step dependencies during annotation, we ensure the assumed causal constraints. Varying experiences and biases act as background factors. If an assessment task is subjective, background factors induce high variability in the answers. To measure the level of noise, each step and input is annotated redundantly. Using this data, we estimate the structural equations by training machine learning models on the input-output data per step.
Step 3. Analysis and Automation: Ignoring training error, the resulting SCM from Step 2 perfectly models the human reasoning process for the given problem. By formulating causal queries on the SCM, we simulate counterfactuals and interventions to describe how inputs and intermediate steps influence the final verdict. Additionally, the workflow serves as a backbone for assessing new documents, either automatically or through human-AI collaboration. Since the workflow decomposes the assessment task into individual steps, various reasoning and assistance scenarios can be tested. In Section 5 we demonstrate that LLMs integrate naturally with the causal graph to solve STRICTA problems reconstructing human reasoning from partial information by turning the causal graph into an LLM program Dohan et al. (2022).
4 Paper Assessment as STRICTA
Figure 2: Paper assessment workflow (excerpt). Each node is associated with a task, description, example, and output format. "…" means there exists a path.
Paper assessment is a critical step in scholarly peer review and can substantially benefit from AI assistance Kuznetsov et al. (2024). We apply the proposed STRICTA framework to biomedical paper assessment. In this task, a reviewer is given the paper and arrives at an overall quality verdict through a series of unobserved reasoning steps. The output of this task is not a full peer review report but the "thought process" underlying it. This framing models the paper assessment process as supported by our expert interviews (see below) and establishes a setting for human-AI collaboration with full human agency over the process. We focus on biomedical paper assessment because of two main reasons: first, our expert interviews below show that paper assessment in biomedicine follows relatively consistent patterns across sub-communities, increasing the flexibility and scope of the resulting assessment workflow. Second, most studies in NLP for peer review focus exclusively on machine learning and NLP domains Staudinger et al. (2024); we contribute to bridging the gap towards other less-represented scientific domains Kuznetsov et al. (2024). Following Definition 1, a workflow for paper assessment describes the mental model of what constitutes a high-quality paper, connecting observable text to the final verdict. We construct the workflow, i.e,. the SCM, reflecting this mental model, populate it through a human annotation study, and analyze human reasoning patterns following the STRICTA framework.